# A Homothetic Gauge Theory

### v2

@@ A.1 Conformal Cartan Geometry
Gauge fixing to the block–homothetic subgroup.
From now on we freeze a future-directed timelike (equivalently : a null) ray in $\mathbb{R}^{4,2}$. On the base $M$ that choice projects to the fixed vector $X_{d}\!\in\!TM$ introduced in Section 2. The remaining structure group is the semidirect product $H=\mathrm{SO}(3,1)\ltimes\mathbb{R}$; its $\mathbb{R}$ factor acts by the block-upper-triangular matrices
@@
Associated bundles and differential forms.
Because every representation of $P$ (and hence of $H\subset P$) yields an associated bundle, all fields used in the main text acquire a natural Cartan meaning:
@@ Appendix B Finding $\lambda$ for imposition of boundary conditions
Conclusion.
The would–be “zero” of $\Omega=e^{-\lambda}$ at any $x_{0}$ inside the support of $\delta_{n}$ is therefore a coordinate artefact: all curvature scalars are finite and causal geodesics can be extended through $x_{0}$. The conformal factor regularises the monopole–like $1/r$ behavior without introducing a physical singularity.
@@ Appendix C Principal $\mathcal{H}$-bundle formulation and curvature
Local expression.
Choosing local sections $\sigma_{\alpha}\colon U_{\alpha}\to P$ we obtain the connection $1$-forms

### v3

@@ A.1 Conformal Cartan Geometry
Gauge fixing to the block–homothetic subgroup. From now on we freeze a future-directed timelike (equivalently : a null) ray in $\mathbb{R}^{4,2}$. On the base $M$ that choice projects to the fixed vector $X_{d}\!\in\!TM$ introduced in Section 2. The remaining structure group is the semidirect product $H=\mathrm{SO}(3,1)\ltimes\mathbb{R}$; its $\mathbb{R}$ factor acts by the block-upper-triangular matrices
@@
Associated bundles and differential forms. Because every representation of $P$ (and hence of $H\subset P$) yields an associated bundle, all fields used in the main text acquire a natural Cartan meaning:
  • 1.
@@ Appendix B Finding $\lambda$ for imposition of boundary conditions
Conclusion. The would–be “zero” of $\Omega=e^{-\lambda}$ at any $x_{0}$ inside the support of $\delta_{n}$ is therefore a coordinate artefact: all curvature scalars are finite and causal geodesics can be extended through $x_{0}$. The conformal factor regularises the monopole–like $1/r$ behavior without introducing a physical singularity.
@@ Appendix C Principal $\mathcal{H}$-bundle formulation and curvature
Local expression. Choosing local sections $\sigma_{\alpha}\colon U_{\alpha}\to P$ we obtain the connection $1$-forms